# Local Policies Enable Zero-shot Long-horizon Manipulation

### v1

@@ I Introduction
The core bottlenecks are: 1) accurately simulating contact dynamics [21] - for which strategies such as domain randomization [17, 22], SDF contacts [23, 14, 15], and real world corrections [16] have shown promise, 2) generating all possible scene and task configurations to ensure trained policies generalize and 3) acquiring long-horizon behaviors themselves, which may require potentially intractable amounts of data for as the horizon grows.
@@ II Related Work
Sim2real approaches in robotics Transfer of RL policies trained with procedural scene generation has produced generalist robot policies for locomotion [8, 9, 11, 10, 12]. However, the robot is often trained for a single skill, such as walking, or a limited set of similar skills, such as walking at different velocities or headings. Sim2real transfer has also been explored for transferring dexterous manipulation skills [17, 22, 18, 47, 19] and contact-rich manipulation [14, 15, 16]. In our work, we train a variety of skills for manipulation and demonstrate zero-shot capabilities on a large set of unseen tasks. We outperform methods that use end-to-end sim2real transfer [48] as well as real world corrections [16], ManipGen is orthogonal to human correction approaches, and can benefit from real-world data as well.
@@ III-A Framework
Any policy class can be used to define the skills, denoted as $p_{\phi_{k}}(a^{t}|g_{k},O^{t})$, which take in the kth sub-goal $g_{k}$ and current observation $O^{t}$. However, note that many manipulation skills (e.g. picking, pushing, turning, etc.) can be decomposed into a policy $\pi_{reach}$ to achieve target poses near objects, $X_{targ,k}$, followed by policy $\pi_{loc}$ for contact-rich interaction. Accordingly, $p_{\phi_{k}}(a^{t}|g_{k},O^{t})=\pi_{reach}(\tau_{reach}|g_{k},O^{t})\pi_{loc}(a_{loc}^{t}|O_{loc}^{t})$. To implement $\pi_{reach}$, we need to interpret language sub-goals $g_{k}$ to take the robot from its current configuration $q_{k,i}$ to some target configuration $q_{k,f}$ such that $X_{ee}$ (the end-effector pose) is close to $X_{targ,k}$. Thus, we structure the VLM’s sub-goal predictions, $g_{k}$, as tuples containing the following information (object, skill). We then interpret these plans into robot poses by pairing any language conditioned pose estimator or affordance model (to predict $X_{targ,k}$) with an inverse kinematics routine (to compute $q_{k,f}$). Motion planning is used to implement $\pi_{reach}$ by predicting trajectories $\tau_{reach}$ to achieve the target configuration $q_{k,f}$ while avoiding collisions.
@@ III-B Training Local Policies for Sim2Real Manipulation
To train local policies, we adapt the standard two-phase training approach [47, 12, 11, 50, 19, 16] in which we first train state-based expert policies using RL, then distill them into visuomotor policies for transfer. Although local policies can generalize automatically across scene arrangements, robot configurations, and object poses, they must be trained across a wide array of objects to foster object-level generalization. To do so, we train a vast array of single-object state-based RL policies and then distill them into generalist visuomotor policies per skill.
@@
Rewards We train RL policies ($\pi_{loc_{k}}$) in simulation using reward functions we design to elicit the desired behavior per skill $k$. We propose a reward framework that encompasses our local skills: $\mathbf{r}=c_{1}r_{ee}+c_{2}r_{obj}+c_{3}r_{ee,obj}+c_{4}r_{action}+c_{5}r_{succ}$. $\mathbf{r}$ specifies behavior for a broad range of manipulation tasks which involve moving the end-effector to specific poses (often right before contact) as well as a target object to desired poses and need to do so while maintaining certain constraints on the relative motion between the end-effector and the object as well as pruning out undesirable actions. $r_{ee}$ encourages reaching/maintaining specific end-effector poses, $r_{obj}$ restricts/encourages specific object poses or joint configurations, $r_{ee,obj}$ constrains the end-effector motion relative to the object(s) in the scene, $r_{action}$ restricts or penalizes undesirable actions and $r_{succ}$ is a binary success reward. Please see the website for detailed descriptions of the task specific reward functions.
@@ III-C Generalist Policies via Distillation
Observation Space Design for Locality For local policies to transfer effectively to the real robot, the observation space and augmentations must be designed with transfer in mind. To imitate a privileged expert, our observation space must be expressive - providing as much information as possible to the agent. The observations must also be local to enable all of the properties of locality, and augmentations must ensure the policy is robust to noisy real world vision.
@@ III-D Zero-shot Text Conditioned Manipulation
For predicting $\tau_{reach}$, while any motion planner can be used, we select Neural MP [25] due to its fast planning time (2s) and strong real-world planning performance. Given $X_{targ,k}$, we compute target joint state $q_{k,f}$, plan with Neural MP open-loop and execute the predicted $\tau_{reach}$ on the robot using a PID joint controller. We then execute the appropriate local policy (as predicted by the VLM) on the robot to perform manipulation. We alternate between motion planning and deploying local policies until the task is complete. Finally, we note that the particular choice of models is orthogonal to our method. For additional details regarding all aspects of our method, please see the Appendix.
@@ IV-C Real World Evaluation
Comparisons We evaluate SOTA text-conditioned manipulation approaches: SayCan [36], LLMTrajGen [45] and VoxPoser [41]. For SayCan, we use our VLM and motion planning system with engineered primitives for interaction; testing the importance of training local policies. We additionally compare against a pre-trained model for manipulation, OpenVLA [77]. For each task, we collect 25 demonstrations on held out objects in held out poses and scene configurations and fine-tune OpenVLA per task. We pass in a text prompt specifying the task, recording the task success rate and number of stages completed.
@@
Across all 5 tasks (Tab. III), we find that ManipGen outperforms all methods, achieving $76\%$ zero-shot success rate overall. Note that we have not trained our local policies on any of these specific objects or in these specific configurations; there is no adaptation in the real world. ManipGen is able to avoid obstacles while performing manipulation of unseen objects in arbitrary poses and configurations. Failure cases for our method resulted from 1) vision failures as open-set detection models such as Grounding Dino [59] detected the wrong object, 2) imperfect motion planning, resulting in collisions with the environment during execution which dropped objects sometimes and 3) local policies failing to manipulate from sub-optimal initial poses. In general, DrawerStore and Tidy are the most challenging tasks due to their horizon, and consequently all methods, including our own perform worse (60% for ours, 20% for best baseline).
@@
SayCan is the strongest baseline (40% success), achieving non-zero success on every task by leveraging the generalization capabilities of vision-language foundation models in a structured manner. However, when initial poses are not ideal or the task requires contact-rich control, pre-defined primitives fall apart (10-20% success). LLMTrajGen, while capable of performing top-down unconstrained pick and place (Cook: $70\%$), only makes partial progress on tasks requiring obstacle avoidance (Replace) or articulated object manipulation (Store) as its prompts struggle to cover those cases well. VoxPoser achieves similar performance on top-down unconstrained pick and place (Cook); however, it struggles to achieve desired rotation on articulated objects and in tight spaces (Replace and Store). Moreover, VoxPoser frequently generates incorrect plans in long-horizon tasks (Tidy). Finally, OpenVLA failed to solve any task, failing to generalize to held out objects and poses even though it was the only method that was given few-shot data. We attempted to evaluate it on its training objects and it still performs poorly with strong pose randomization. For additional details regarding evaluation protocol, experiment conditions and deployment, please see the Appendix.
V Discussion and Limitations
@@
We present ManipGen, a method for solving long-horizon manipulation tasks with unseen objects in unseen configurations by training generalist policies for sim2real transfer. We propose local policies, a novel policy class for sim2real transfer that is pose, skill order and scene configuration invariant, enabling it generalize broadly. For deployment, we take advantage of the broad generalization capabilities of foundation models for vision, language and motion planning to solve long-horizon manipulation tasks from text prompts. Across 50 real-world long-horizon manipulation tasks, our method achieves 76% zero-shot success, outperforming SOTA planning and imitation methods on every task.
ManipGen also leaves significant room for future work, for which we briefly describe limitations and outline potential solutions. First, as a modular framework, ManipGen is susceptible to cascading errors from any of the modules, particularly from the high-level planner providing incorrect plans and the pose-estimation latching onto the wrong object. Training value functions per module, incorporating closed-loop re-planning and double checking as well as online-adaptation of the modules (as they are all learnable) are all potential techniques that could be applied to diminish the impact of cascading errors. Next, due to the dependence on depth sensing, for pose-estimation, motion planning and local policy execution, ManipGen performs struggles with shiny, reflective and transparent objects. To that end, methods that learn depth from RGB [78, 79] could directly resolve this limitation. Finally, as interaction is learned in simulation, ManipGen is limited to learning behaviors that can be accurately simulated, limiting its direct applicability to tasks with state-changes and complex contacts. We note that the local policy framework described in this work, while uniquely suited for sim2real transfer, is amenable to training on real-world data as well.
@@ VI-B Rewards
Pick involves moving the gripper so the object can be easily grasped. Instead of encouraging the agent to move towards the overall object pose, which is not necessarily the pose to achieve for grasping, we provide dense signal for learning to grasp using pre-sampled grasp poses ($\{X_{targ}\}$). We encourage the agent to minimize the key-point distance [80] to the nearest grasp pose:

### v2

@@ I Introduction
The core bottlenecks are: 1) accurately simulating contact dynamics [21] – for which strategies such as domain randomization [17, 22], SDF contacts [23, 14, 15], and real world corrections [16] have shown promise, 2) generating all possible scene and task configurations to ensure trained policies generalize and 3) acquiring long-horizon behaviors themselves, which may require potentially intractable amounts of data for as the horizon grows.
@@ II Related Work
Sim2real approaches in robotics Transferring RL policies trained with procedural scene generation enables generalist robot locomotion [8, 9, 11, 10, 12]. However, these policies are typically limited to a single skill, such as walking, or slight variations like different velocities or headings. Sim2real transfer has also been explored for dexterous manipulation skills [17, 22, 18, 47, 19] and contact-rich manipulation [14, 15, 16]. In our work, we train a variety of skills for manipulation and demonstrate zero-shot generalization on numerous unseen tasks. We outperform methods that use end-to-end sim2real transfer [48] as well as real world corrections [16]. ManipGen is orthogonal to human correction approaches, and can benefit from real-world data as well.
@@ III-A Framework
Any policy class can be used to define the skills, denoted as $p_{\phi_{k}}(a^{t}|g_{k},O^{t})$, which take in the $k$th sub-goal $g_{k}$ and current observation $O^{t}$. However, note that many manipulation skills (e.g. picking, pushing, turning, etc.) can be decomposed into a policy $\pi_{reach}$ to achieve target poses near objects, $X_{targ,k}$, followed by policy $\pi_{loc}$ for contact-rich interaction. Accordingly, $p_{\phi_{k}}(a^{t}|g_{k},O^{t})=\pi_{reach}(\tau_{reach}|g_{k},O^{t})\pi_{loc}(a_{loc}^{t}|O_{loc}^{t})$. To implement $\pi_{reach}$, we need to interpret language sub-goals $g_{k}$ to take the robot from its current configuration $q_{k,i}$ to some target configuration $q_{k,f}$ such that $X_{ee}$ (the end-effector pose) is close to $X_{targ,k}$. We structure the VLM’s sub-goal predictions $g_{k}$ as (object, skill) tuples and interpret these plans into robot poses by pairing any language conditioned pose estimator or affordance model (to predict $X_{targ,k}$) with an inverse kinematics routine (to compute $q_{k,f}$). Motion planning is used to implement $\pi_{reach}$ by predicting collision-avoidant trajectories $\tau_{reach}$ to achieve the target configuration $q_{k,f}$.
@@ III-B Training Local Policies for Sim2Real Manipulation
To train local policies, we adapt the standard two-phase training approach [47, 12, 11, 50, 19, 16] in which we first train state-based expert policies using RL, then distill them into visuomotor policies for transfer. While local policies can generalize across scene arrangements, robot configurations, and object poses, broad object-level generalization requires diverse training data. To achieve this, we train a vast array of single-object state-based RL policies and distill them into generalist visuomotor policies per skill.
@@
Rewards We train RL policies ($\pi_{loc_{k}}$) in simulation using reward functions we design to elicit the desired behavior per skill $k$. We propose a reward framework that encompasses our local skills: $\mathbf{r}=c_{1}r_{ee}+c_{2}r_{obj}+c_{3}r_{ee,obj}+c_{4}r_{action}+c_{5}r_{succ}$. $\mathbf{r}$ specifies behavior for a broad range of manipulation tasks which involve moving the end-effector to specific poses (often right before contact) as well as a target object to desired poses and need to do so while maintaining certain constraints on the relative motion between the end-effector and the object as well as pruning out undesirable actions. $r_{ee}$ encourages reaching/maintaining specific end-effector poses, $r_{obj}$ restricts/encourages specific object poses or joint configurations, $r_{ee,obj}$ constrains the end-effector motion relative to the object(s) in the scene, $r_{action}$ restricts or penalizes undesirable actions and $r_{succ}$ is a binary success reward.
@@ III-C Generalist Policies via Distillation
Observation Space Design for Locality For local policies to transfer effectively to the real robot, the observation space and augmentations must be designed with transfer in mind. To imitate a privileged expert, our observation space must be expressive – providing as much information as possible to the agent. The observations must also be local to enable all of the properties of locality, and augmentations must ensure the policy is robust to noisy real world vision.
@@ III-D Zero-shot Text Conditioned Manipulation
For predicting $\tau_{reach}$, while any motion planner can be used, we select Neural MP [25] for its fast (2s) and strong real-world planning. Given $X_{targ,k}$, we compute target joint state $q_{k,f}$, plan with Neural MP open-loop and execute the predicted $\tau_{reach}$ on the robot using a PID joint controller. We then execute the appropriate local policy (as predicted by the VLM) on the robot to perform manipulation. We alternate between motion planning and deploying local policies until the task is complete. Finally, we note that the particular choice of models is orthogonal to our method.
@@ IV-C Real World Evaluation
Comparisons We evaluate SOTA text-conditioned manipulation approaches: SayCan [36], LLMTrajGen [45] and VoxPoser [41]. For SayCan, we use our VLM and motion planning system with engineered interaction primitives to assess the role of training local policies. We additionally compare against a pre-trained manipulation model OpenVLA [77], fine-tuning it per task using 25 demonstrations on held-out objects, poses, and scene configurations. Given a text prompt, we record task success rates and stages completed.
@@
Across all 5 tasks (Tab. III), ManipGen achieves $76\%$ zero-shot success rate, outperforming all methods. Notably, our local policies are not trained on any of these specific objects or in these specific configurations and require no real-world adaptation. ManipGen is able to avoid obstacles while performing manipulation of unseen objects in arbitrary poses and configurations. Failure cases stemmed from 1) vision failures as open-set detection models such as Grounding Dino [59] detected the wrong object, 2) imperfect motion planning, resulting in collisions with the environment during execution which dropped objects sometimes and 3) local policies failing to manipulate from sub-optimal initial poses. In general, DrawerStore and Tidy are the most challenging tasks due to their horizon, and consequently all methods, including our own perform worse (60% for ours, 20% for best baseline).
@@
SayCan is the strongest baseline (40% success), achieving non-zero success on every task by leveraging the generalization capabilities of vision-language foundation models in a structured manner. However, when initial poses are not ideal or the task requires contact-rich control, pre-defined primitives fall apart (10-20% success). LLMTrajGen excelled at top-down pick-and-place (Cook: 70%) but struggled with obstacle avoidance (Replace) and articulated object manipulation (Store) due to limited prompt coverage. VoxPoser matches LLMTrajGen in Cook but failed at precise rotations and tight-space tasks (Replace, Store) and frequently generated incorrect plans for long-horizon tasks (Tidy). Finally, OpenVLA failed to solve any task, unable to generalize to held-out objects and poses even with few-shot data. We attempted to evaluate it on its training objects and it still performs poorly with strong pose randomization.
V Discussion
@@
We present ManipGen, a method for long-horizon manipulation tasks with unseen objects and configurations via generalist policies for sim2real transfer. We propose local policies, a novel policy class invariant to pose, skill order, and scene configuration, enabling broad generalization. For deployment, we leverage foundation models for vision, language and motion planning to solve long-horizon manipulation tasks from text prompts. Across 50 real-world long-horizon manipulation tasks, ManipGen achieves 76% zero-shot success, surpassing SOTA planning and imitation methods.
@@ VI-B Rewards
Pick involves moving the gripper so the object can be easily grasped. Instead of encouraging the agent to move towards the overall object pose, which is not necessarily the pose to achieve for grasping, we provide dense signal for learning to grasp using pre-sampled grasp poses ($\{X_{targ}\}$). We encourage the agent to minimize the key-point distance [allshire2022transferring] to the nearest grasp pose: